What is the maximum length of a closed pipe that will produce just audible sound(velocity of sound=336m/s)?

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
Minimum audible frequency = 20 Hz
 
Now
 
v /4 * L = 20
L = 336/80 = 4.2 m
 
Regards
Arun (askIITians forum expert)